# Bayesian inversion with Student’s $t$ priors based on Gaussian scale mixtures

### v1

@@ 5.2 Image deblurring
In image deblurring, we aim at recovering the original sharp image from the corresponding noisy blurred image using a mathematical model of the blurring process.
If we denote the image domain by $\mathcal{D}=([0,1]\times[0,1])$, and $\bm{t}=[t_{1},t_{2}]^{\mathsf{T}}\in\mathcal{D}$, $\bm{s}=[s_{1},s_{2}]^{\mathsf{T}}\in\mathcal{D}$ then the blurred image $y$ can be obtained from the true image $x$ using the Fredholm integral operator as follows Lu et al. (2010)
@@
We estimate posterior under Student’s $t$ prior (with $\nu\sim\text{Ga}_{\nu>1}(\nu;\,2,0.1)$) using the Gibbs sampler. Next, we compare results with those obtained by NUTS with Laplace MRF prior that is defined in two-dimensional case as follows
$p(\bm{x})=\left(\frac{\delta}{2}\right)^{d}\exp\left[-\delta\left(\|\bm{D}^{(1)}\bm{x}\|_{1}+\|\bm{D}^{(2)}\bm{x}\|_{1}\right)\right],$ (34)
where difference matrices $\bm{D}^{(1)}$ and $\bm{D}^{(2)}$ are horizontal and vertical difference matrices, parameter $\delta\in\mathbb{R}_{>0}$ is the inverse scale, and $\left\|\cdot\right\|_{1}$ denotes the $\ell_{1}$-norm.
The input parameters to the Gibbs and NUTS samples are the same as in one-dimensional deconvolution example. In Figure 9, we show posterior mean and logarithm of standard deviation for two deblurred images obtained with Student’s $t$ prior and Laplace prior. We plot the logarithm of standard deviation to facilitate the interpretation for the Student’s $t$ prior, as maximum values of standard deviation occur at single pixels of the image. We note for Student’s $t$ prior the minimum value of the standard deviation is $1.286\times 10^{-4}$ and the maximum value is $0.041$ (if we omit two points exhibiting larger standard deviation values of $0.639$ and $0.675$). Whereas, for Laplace prior the standard deviation ranges from $0.128$ to $0.259$.
The relative reconstruction error is $0.218$ for Student’s $t$ and $0.215$ for Laplace. Despite both priors yield comparable relative errors, we can clearly see that the mean of posterior based on Student’s $t$ prior is sharper, contains less noise and has smaller posterior uncertainty. It can be also noted that for the Laplace prior the larger values of standard deviation are located at the edges pointing to the higher uncertainty in the boundary regions.
Figure 9: Comparison of posterior statistics for the target parameter $\bm{x}$ using the Student’s $t$ prior (top row) and the Laplace prior (bottom row). Posterior mean with relative error $\varepsilon_{\mathrm{rel}}$ (left) and logarithm of the standard deviation (right).
In addition, for Student’s $t$ , we report posterior statistics of the local scale parameter $\bm{w}$ in Figure 10, and chain statistics of scalar parameters (degrees of freedom $\nu$ and the global scale $\tau$) in Figure 11. In the Laplace prior case, the posterior mean and standard deviation of the inverse scale parameter $\delta$ are $3.214$ and $0.151$, respectively.
@@
Figure 11: Posterior statistics for $\nu$ (top row) and $\tau$ (bottom row), obtained under the Student’s $t$ prior using Gibbs: histogram, chain, cumulative mean and sample autocorrelation.
@@ 6 Summary and conclusions
We developed a computational framework for solving linear Bayesian inverse problems in which the unknown can be either discontinuous or smooth. Our method involves flexible priors based on Markov random fields combined with Student’s $t$ distribution. We solve a hierarchical Bayesian inverse problem that allows joint estimation of the degrees of freedom parameter of the Student’s $t$ distribution. Flexible estimation of degrees of freedom provides an opportunity to tailor the prior for recovering not only sharp but also smooth features depending on the task.
Due to the hierarchical structure and the Student’s $t$ distributions involved in the prior, sampling of the posterior distribution is challenging. To address the issue, we employed a Gaussian scale mixture representation of the Student’s $t$ distribution, that allowed us to employ an efficient Gibbs sampler. We validated numerically our suggested Gibbs sampler by comparing its results to the solution obtained via NUTS.
In the numerical experiments, we showed that Student’s $t$ hierarchical prior can be applied for recovering blocky objects as well as the objects containing smooth features. To access the performance of the proposed method in Bayesian inversion, we compared Student’s $t$ priors with more traditional prior models based on Laplace Markov random fields commonly used for promoting sharp features. As compared to Laplace priors, Student’s $t$ priors enable the computation of sharp point estimates of the posterior while reducing the posterior uncertainty.
As a future work, we consider employing a higher-order difference matrix in the MRF structure to add more structural dependencies. This could be potentially beneficial for the smooth case scenario as opposed to our current implementation based on the first-order difference matrix, which models connections in the two neighbouring pixels only. In addition, we would like to extend the method to the higher dimensional inverse problems. This essentially requires application of efficient methods to sample from high-dimensional Gaussian distributions.
@@ Acknowledgments
This work was supported by the Research Council of Finland through the Flagship of Advanced Mathematics for Sensing, Imaging and Modelling, and the Centre of Excellence of Inverse Modelling and Imaging (decision numbers 359183 and 353095, respectively). In addition, AS was supported by a doctoral studies research grant from the Vilho, Yrjö and Kalle Väisälä Foundation of the Finnish Academy of Science and Letters.

### v2

@@ 5.2 Image deblurring
In image deblurring, we aim to recover the original sharp image from the corresponding noisy blurred image using a mathematical model of the blurring process. If we denote the image domain by $\mathcal{D}=([0,1]\times[0,1])$, and $\bm{t}=[t_{1},t_{2}]^{\mathsf{T}}\in\mathcal{D}$, $\bm{s}=[s_{1},s_{2}]^{\mathsf{T}}\in\mathcal{D}$ then the blurred image $y$ can be obtained from the true image $x$ using the Fredholm integral operator as follows [58]
@@
We estimate posterior under the Student’s $t$ prior (with $\nu\sim\text{Ga}_{\nu>1}(\nu;\,2,0.1)$) using the Gibbs sampler. Next, we compare results with those obtained by NUTS with Laplace MRF prior. This prior is defined in the two-dimensional case as
$p(\bm{x})=\left(\frac{1}{2\tau}\right)^{d}\exp\left[-\frac{1}{\tau}\left(\|\bm{D}^{(1)}\bm{x}\|_{1}+\|\bm{D}^{(2)}\bm{x}\|_{1}\right)\right],$ (32)
where $\bm{D}^{(1)}$ and $\bm{D}^{(2)}$ are horizontal and vertical difference matrices, parameter $\tau\in\mathbb{R}_{>0}$ is the scale, and $\left\|\cdot\right\|_{1}$ denotes the $\ell_{1}$-norm.
The input parameters to the Gibbs and NUTS methods are the same as in the one-dimensional deconvolution example. In Figure 9, we show the posterior mean and logarithm of standard deviation for two deblurred images obtained with Student’s $t$ prior and Laplace prior. We plot the logarithm of standard deviation to facilitate the interpretation for the Student’s $t$ prior, as maximum values of standard deviation occur at single pixels of the image. We note for the Student’s $t$ prior, the minimum value of the standard deviation is $1.286\times 10^{-4}$ and the maximum value is $0.041$ (if we omit two points exhibiting larger standard deviation values of $0.639$ and $0.675$); whereas for Laplace prior, the standard deviation ranges from $0.128$ to $0.259$.
The relative reconstruction error is $0.218$ for the $t$ prior and $0.215$ for the Laplace prior. Despite both priors yield comparable relative errors, we can clearly see that the mean of posterior based on the Student’s $t$ prior is sharper, contains less noise, and has smaller posterior uncertainty. It can be also noted that for the Laplace prior the larger values of standard deviation are located at the edges pointing to the higher uncertainty in the boundary regions.
Figure 9: Comparison of posterior statistics for the target parameter $\bm{x}$ using the Student’s $t$ prior (top row) and the Laplace prior (bottom row). Posterior mean with relative error $\varepsilon_{\mathrm{rel}}$ (left) and the logarithm of the standard deviation (right).
In addition, for the Student’s $t$ prior, we report posterior statistics of the local scale parameter $\bm{w}$ in Figure 10, and chain statistics of scalar parameters (degrees of freedom $\nu$ and the global scale $\tau$) in Figure 11. In the Laplace prior case, the posterior mean and standard deviation of the scale parameter $\tau$ are $0.312$ and $0.015$, respectively.
@@
Figure 11: Posterior statistics for $\nu$ (top row) and $\tau$ (bottom row), obtained under the Student’s $t$ prior using Gibbs: histogram, chain, cumulative mean, and sample autocorrelation.
@@ 6 Summary and conclusions
We developed a computational framework for solving linear Bayesian inverse problems, in which the unknown can be either discontinuous or smooth. Our method involves flexible priors based on Markov random fields combined with Student’s $t$ distribution. We solve a hierarchical Bayesian inverse problem that allows joint estimation of the degrees of freedom parameter of Student’s $t$ distribution. The estimation of the degrees of freedom tailors the prior for recovering not only sharp but also smooth features depending on the underlying unknown solution. Modeling the degrees of freedom as a hyperparameter, however, adds an extra layer to the hierarchical model, which increases the inference complexity. At the same time, this does not affect the computational complexity, as the parameter is one-dimensional. Moreover, we note that the automatic inference of the degrees of freedom parameter eliminates the need for expensive manual tuning of this parameter.
Due to the hierarchical structure and Student’s $t$ distribution involved in the prior, sampling of the posterior distribution is challenging. To address the issue, we employed a Gaussian scale mixture representation of Student’s $t$ distribution, that allowed us to employ an efficient Gibbs sampler. We validated numerically our suggested Gibbs sampler by comparing its results to the solution obtained via NUTS.
In the numerical experiments, we showed that Student’s $t$ hierarchical prior can be applied for recovering blocky objects as well as objects containing smooth features. To access the performance of the proposed method in Bayesian inversion, we compared Student’s $t$ priors with more traditional prior models based on Laplace and Cauchy Markov random fields commonly used for promoting sharp features. As compared to Laplace priors, Student’s $t$ priors enable the computation of sharp point estimates of the posterior while reducing the posterior uncertainty. At the same time, the $t$ priors are flexible enough to capture smooth features in the solution, whereas the Cauchy MRF priors can only promote sharp behavior. We also observed that when the unknown signals are sharp, our model benefits from weak prior information (i.e., no truncation of the degrees of freedom hyperprior), while for smooth targets the prior information is preferably strong (i.e., truncation of the degrees of freedom hyperprior).
As a future work, we consider employing a higher-order difference matrix in the MRF structure to add more structural dependencies. This could be potentially beneficial for the smooth case scenario as opposed to our current implementation based on the first-order difference matrix, which models connections in the two neighboring pixels only. In addition, we would like to extend the method to the higher dimensional inverse problems. This essentially requires the application of efficient methods to sample from high-dimensional Gaussian distributions.
@@ Acknowledgments
This work was supported by the Research Council of Finland through the Flagship of Advanced Mathematics for Sensing, Imaging and Modelling, and the Centre of Excellence of Inverse Modelling and Imaging (decision numbers 359183 and 353095, respectively). In addition, AS was supported by the doctoral studies research grant from the Vilho, Yrjö and Kalle Väisälä Foundation of the Finnish Academy of Science and Letters.